# Reciprocity Theorem and Fundamental Transfer Matrix

### v1

@@ 1 Introduction
  • Theorem 1 (Reciprocity in 1D) Let $v:\mathbb{R}\to\mathbb{C}$ be a short-range potential. Then its left and right transmission amplitudes coincide.
@@
  • Theorem 2 (Reciprocity in 2D and 3D) Let $d\in\{1,2\}$ and $v:\mathbb{R}^{d+1}\to\mathbb{C}$ be a short-range potential. Then the scattering amplitude $\mathfrak{f}({\mathbf{n}}_{0},{\mathbf{n}})$ of $v$ satisfies
@@ 2 Dynamical formulation of stationary scattering in 2D and 3D
Next, we introduce a useful notation: For each $\mathbf{u}:=({\rm u}_{x},{\rm u}_{y},{\rm u}_{z})\in\mathbb{R}^{3}$, we use $\vec{\rm u}$ to denote $({\rm u}_{y},{\rm u}_{z})$. This allows us to identify $\mathbf{u}$ with $({\rm u}_{x},\vec{\rm u})$. In particular, $\vec{\rm r}:=(y,z)$ and $\mathbf{r}=(x,\vec{\rm r})$. We refer to $\vec{\rm u}$ as the projection of $\mathbf{u}$ onto the $y$-$z$ plane.
The above discussion applies to the scattering problem defined by the Schrödinger equation (20) in 2D once we set the $z$ component of all the relevant vectorial quantities to zero and neglect them in our calculations. In particular, we have $\vec{\rm r}=y$.
Because $v$ is a short-range potential, solutions $\psi(x,\vec{\rm r})$ of (20) tend to superpositions of plane waves as $x\to\pm\infty$. This means that they admit asymptotic expressions of the form
@@
$\mathscr{D}_{k}:=\left\{\>{\vec{p}}\in\mathbb{R}^{2}\,|\,{\vec{p}}^{\,2}<k^{2}\,\right\},$ (24)
$\varpi({\vec{p}}):=\left\{\begin{array}[]{ccc}\sqrt{k^{2}-{\vec{p}}^{\,2}}&{\mbox{\rm for}}&{\vec{p}}\in\mathscr{D}_{k},\\ i\sqrt{{\vec{p}}^{\,2}-k^{2}}&{\mbox{\rm for}}&{\vec{p}}\notin\mathscr{D}_{k},\end{array}\right$. (27)
@@
According to (43) – (52), $\widehat{\mathscr{V}}(x)$, $\widehat{\mathbf{H}}(x)$, and consequently $\widehat{\mathbf{M}}$ and its entries $\widehat{M}_{ij}$ are linear integral operators. Reference [17] offers various examples where the latter can be computed analytically and used to obtain the exact solution of the corresponding scattering problems.
@@ 3 Statement of the reciprocity relation in terms of $\widehat{\mathbf{M}}$
Equations (64) – (67), (74), and (75) allow us to identify the reciprocity relations (68) – (71) with
@@
where $\hat{O}:L^{2}(\mathscr{D}_{k})\to L^{2}(\mathscr{D}_{k})$ is any densely-define linear operator, and $\hat{O}^{\dagger}$ denotes its adjoint. Equations (84) allow us to express the reciprocity relations (76) – (78) as the following conditions on the entries of the fundamental transfer matrix.
@@
Because $(\widehat{\mathcal{P}}\widehat{\mathcal{T}})^{-1}=\widehat{\mathcal{P}}\widehat{\mathcal{T}}$, (85) and (86) state that $\widehat{M}_{22}^{-1}\widehat{M}_{21}\widehat{\varpi}$ and $\widehat{M}_{12}\widehat{M}_{22}^{-1}\widehat{\varpi}$ are $\widehat{\mathcal{P}}\widehat{\mathcal{T}}-$anti-pseudo-Hermitian [14, 15]. Notice also that, because $\widehat{\varpi}$ and $\widehat{\varpi}^{-1}$ act in $L^{2}(\mathscr{D}_{k})$ as linear Hermitian operators, we can express (85) – (87) in the form
@@ 4 The property of $\widehat{\mathbf{M}}$ that is responsible for reciprocity
    $\widehat{\boldsymbol{\Omega}}\,\widehat{\mathcal{P}}\widehat{\mathcal{T}}\,\widehat{\varpi}^{-1}\widehat{\mathbf{H}}(x)\,(\widehat{\boldsymbol{\Omega}}\,\widehat{\mathcal{P}}\widehat{\mathcal{T}}\widehat{\varpi}^{-1})^{-1}=-\widehat{\mathbf{H}}(x)^{\dagger},$ (109)

### v2

@@ 1 Introduction
  • Theorem 1 (Reciprocity in 1D) Let $v:\mathbb{R}\to\mathbb{C}$ be a real or complex short-range potential. Then its left and right transmission amplitudes coincide.
@@
  • Theorem 2 (Reciprocity in 2D and 3D) Let $d\in\{1,2\}$ and $v:\mathbb{R}^{d+1}\to\mathbb{C}$ be a real or complex short-range potential. Then the scattering amplitude $\mathfrak{f}({\mathbf{n}}_{0},{\mathbf{n}})$ of $v$ satisfies
@@ 2 Dynamical formulation of stationary scattering in 2D and 3D
Next, we introduce a useful notation: For each $\mathbf{u}:=({\rm u}_{x},{\rm u}_{y},{\rm u}_{z})\in\mathbb{R}^{3}$, we use $\vec{\rm u}$ to denote $({\rm u}_{y},{\rm u}_{z})$. This allows us to identify $\mathbf{u}$ with $({\rm u}_{x},\vec{\rm u})$. We refer to $\vec{\rm u}$ as the projection of $\mathbf{u}$ onto the $y$-$z$ plane.
The above discussion applies to the scattering problem defined by the Schrödinger equation (20) in 2D once we set the $z$ component of all the relevant vectorial quantities to zero and neglect them in our calculations. In particular, for $\mathbf{u}:=({\rm u}_{x},{\rm u}_{y})\in\mathbb{R}^{2}$, we have $\vec{\rm u}:={\rm u}_{y}$ and $\mathbf{u}=(u_{x},\vec{u})$.
Because $v$ is a short-range potential, bounded solutions $\psi(x,\vec{\rm r})$ of (20) tend to superpositions of plane waves as $x\to\pm\infty$. This means that they admit asymptotic expressions of the form
@@
${\vec{\rm r}:=\left\{\>\begin{array}[]{cc}y&\mbox{in 2D},\\ (y,z)&\mbox{in 3D},\end{array}\right.}$ ${\vec{p}:=\left\{\>\begin{array}[]{cc}p_{y}&\mbox{in 2D},\\ (p_{y},p_{z})&\mbox{in 3D},\end{array}\right.}$ (28)
$\mathscr{D}_{k}:=\left\{\>{\vec{p}}\in\mathbb{R}^{2}\,|\,{\vec{p}}^{\,2}<k^{2}\,\right\},$ $\varpi({\vec{p}}):=\left\{\begin{array}[]{ccc}\sqrt{k^{2}-{\vec{p}}^{\,2}}&{\mbox{\rm for}}&{\vec{p}}\in\mathscr{D}_{k},\\ i\sqrt{{\vec{p}}^{\,2}-k^{2}}&{\mbox{\rm for}}&{\vec{p}}\notin\mathscr{D}_{k},\end{array}\right$. (31)
@@
According to (47) – (56), $\widehat{\mathscr{V}}(x)$, $\widehat{\mathbf{H}}(x)$, and consequently $\widehat{\mathbf{M}}$ and its entries $\widehat{M}_{ij}$ are linear integral operators. Reference [17] offers various examples where the latter can be computed analytically and used to obtain the exact solution of the corresponding scattering problems. Among notable applications of the dynamical formulation of stationary scattering, which makes use of the concept of fundamental transfer matrix, are a singularity-free treatment of the delta-function point interactions in 2D and 3D [23] and the discovery of short-range potentials for which the Born approximation is exact. These results extend to electromagnetic scattering [24, 25]. Ref. [26] uses the standard approach to stationary scattering based on the Lippmann-Schwinger equation to offer an independent confirmation of the results on the exactness of the $N$-th order Born approximation that were originally obtained using the fundamental transfer matrix.
We can view the effective Hamiltonian $\widehat{\mathbf{H}}(x)$ as a linear operator acting in the space of square-integrable two-component wave functions $L^{2}(\mathbb{R}^{d})\otimes\mathbb{C}^{2\times 1}$. It is easy to see that in general this operator fails to be Hermitian. This is also true when $v$ is a real potential. In this case the non-Hermiticity of $\widehat{\mathbf{H}}(x)$ does not conflict with the unitarity of the time-evolution of the original system, because the latter is determined by the Schrödinger operator $\nabla^{2}+v(\mathbf{r})$ which is a Hermitian operator acting in the Hilbert space $L^{2}(\mathbb{R}^{d+1})$ of the original system.
@@ 3 Statement of the reciprocity relation in terms of $\widehat{\mathbf{M}}$
As we show in the appendix, Eqs. (68) – (71), (78), and (79) allow us to identify the reciprocity relations (72) – (75) with
@@
where $\hat{O}:L^{2}(\mathscr{D}_{k})\to L^{2}(\mathscr{D}_{k})$ is any densely-define linear operator, and $\hat{O}^{\dagger}$ denotes its adjoint. In the appendix we use (88) to express the reciprocity relations (80) – (82) as the following conditions on the entries of the fundamental transfer matrix.
@@
Because $(\widehat{\mathcal{P}}\widehat{\mathcal{T}})^{-1}=\widehat{\mathcal{P}}\widehat{\mathcal{T}}$, (89) and (90) state that $\widehat{M}_{22}^{-1}\widehat{M}_{21}\widehat{\varpi}$ and $\widehat{M}_{12}\widehat{M}_{22}^{-1}\widehat{\varpi}$ are $\widehat{\mathcal{P}}\widehat{\mathcal{T}}-$anti-pseudo-Hermitian [14, 15]. In the appendix, we use the fact that $\widehat{\varpi}$ and $\widehat{\varpi}^{-1}$ act in $L^{2}(\mathscr{D}_{k})$ as linear Hermitian operators to express (89) – (91) in the form
@@ 6 Concluding remarks
Appendix: Derivation of Eqs. (80) – (82) and (89) – (94)
First, we use (52), (68), and (72) to establish
$\langle\vec{\rm k}|\,\widehat{M}_{22}^{-1}\widehat{M}_{21}\widehat{\varpi}|\vec{\rm k}_{0}\rangle=\langle-\,\vec{\rm k}_{0}|\,\widehat{M}_{22}^{-1}\widehat{M}_{21}\widehat{\varpi}|\!-\vec{\rm k}\rangle.$ (162)
In view of (78), this equation implies (80). Similarly, (70) and (73) imply that (162) holds with $\widehat{M}_{22}^{-1}\widehat{M}_{21}$ changed to $\widehat{M}_{12}\widehat{M}_{22}^{-1}$. This proves (81). Eq. (82) follows from (69), (71), and (74), for we can use them to show that
$\langle\vec{\rm k}|(\widehat{M}_{11}-\widehat{M}_{12}\widehat{M}_{22}^{-1}\widehat{M}_{21})\widehat{\varpi}|\vec{\rm k}_{0}\rangle$ $=$ $\langle-\vec{\rm k}_{0}|\widehat{M}_{22}^{-1}\widehat{\varpi}|-\!\vec{\rm k}\rangle$
$=$ $\langle\vec{\rm k}_{0}|\,\widehat{\mathcal{P}}\widehat{M}_{22}^{-1}\widehat{\varpi}\widehat{\mathcal{P}}\,|\vec{\rm k}\rangle.$
This completes the derivation of (80) – (82).
Next, we use (80) and (88) to show that
$\langle\vec{\rm k}_{0}|\,\widehat{\mathcal{T}}\big(\widehat{M}_{22}^{-1}\widehat{M}_{21}\widehat{\varpi}\big)^{\dagger}\widehat{\mathcal{T}}\,|\vec{\rm k}\rangle$ $=\langle\vec{\rm k}_{0}|\,\widehat{\mathcal{P}}\widehat{M}_{22}^{-1}\widehat{M}_{21}\widehat{\varpi}\widehat{\mathcal{P}}\,|\vec{\rm k}\rangle.$ (163)
Because $\widehat{\mathcal{T}}^{2}=\widehat{I}$ and $\widehat{\mathcal{P}},\widehat{\mathcal{T}}]=\widehat{0}$, (163) implies (89). Equation (90) follows similarly from (81) and (88), as they imply (163) with $\widehat{M}_{22}^{-1}\widehat{M}_{21}$ changed to $\widehat{M}_{12}\widehat{M}_{22}^{-1}$. To derive (91), we note that according to (82) and (88),
$\langle\vec{\rm k}|(\widehat{M}_{11}-\widehat{M}_{12}\widehat{M}_{22}^{-1}\widehat{M}_{21})\,\widehat{\varpi}|\vec{\rm k}_{0}\rangle$ $=\langle\vec{\rm k}|\,\widehat{\mathcal{T}}\big(\widehat{\mathcal{P}}\widehat{M}_{22}^{-1}\widehat{\varpi}\widehat{\mathcal{P}}\big)^{\dagger}\widehat{\mathcal{T}}\,|\vec{\rm k}_{0}\rangle.$ (164)
Because $\widehat{\mathcal{P}}^{\dagger}=\widehat{\mathcal{P}}$ and $(\widehat{\mathcal{P}}\widehat{\mathcal{T}})^{2}=\widehat{I}$, (164) implies (91). This completes the derivation of (89) – (91).
Equations (92) – (94) follow from (89) – (91) if we multiply the left-hand sides of the latter equations by $\widehat{\varpi}^{-1}$ from the left and making use of (95) and the fact that $\widehat{\mathfrak{T}}^{-1}=\widehat{\varpi}\,\widehat{\mathcal{P}}\widehat{\mathcal{T}}$.
It is easy to see that the logical chain of arguments we use to derive (92) – (94) from (72) – (75) can be used in the reverse order to derive (72) – (75) from (92) – (94). This proves the equivalence of the reciprocity relations (72) – (75) and either of the triplets of equations (80) – (82), (89) – (91), and (92) – (94).